# The amplifier effect of artificial agents in social contagion

### v1

@@ Abstract
Recent advances in artificial intelligence have led to the proliferation of artificial agents in social contexts, ranging from education to online social media and financial markets, among many others. The increasing rate at which artificial and human agents interact makes it urgent to understand the consequences of human-machine interactions for the propagation of new ideas, products, and behaviors in society. Across two distinct empirical contexts, we find here that artificial agents lead to significantly faster and wider social contagion. To this end, we replicate a choice experiment previously conducted with human subjects by using artificial agents powered by large language models (LLMs). We use the experiment’s results to measure the adoption thresholds of artificial agents and their impact on the spread of social contagion. We find that artificial agents tend to exhibit lower adoption thresholds than humans, which leads to wider network-based social contagions. Our findings suggest that the increased presence of artificial agents in real-world networks may accelerate behavioral shifts, potentially in unforeseen ways.
The increasing role of artificial agents in human activities is not guaranteed to promote more inclusive, healthy, and sustainable societies [1]. As a result, even before the recent popularization of large language models (LLMs), scholars from multiple disciplines have called for the development of an interdisciplinary science of machine behavior where the behavior of AI agents is studied with similar tools to those traditionally used to understand human behaviors [2]. The proliferation of large language models such as ChatGPT and the increased impact of algorithms on large-scale social dynamics makes it urgent to advance in this direction [1, 3]. However, despite some early efforts [4, 5], the impacts of autonomous agents driven by LLMs (hereafter, artificial agents) on social contagion processes have not yet been widely examined. Here we ask: What role could artificial agents play in the spreading of social contagions? To answer, we consider simple artificial agents obtained by prompting popular LLMs (OpenAI’s gpt-3.5-turbo and Google’s gemini-1.5-flash) with the demographic information and political orientation of the human subjects in a previous study [6]. We perform the first measurement of the artificial agents’ thresholds to adoption, namely, the level of social reinforcement they need before supporting a new policy or adopting a new technology [7]. Our findings indicate that across both contexts, compared to human subjects, artificial agents exhibit a lower threshold to adoption, which invariably results in wider diffusion. The higher the proportion of artificial agents in a human-LLM network, the wider the diffusion.
@@
Here, we aim to advance this debate by providing empirical evidence on the artificial agents’ thresholds. To this end, we adopt two different LLMs – OpenAI’s gpt-3.5-turbo model (hereafter, GPT) and Google’s gemini-flash-1.5 model (hereafter, Gemini) – to simulate the responses of survey participants to the same two choice-based conjoint experiments performed in ref. [6] with human subjects. In conjoint experiments, participants face a set of choice tasks where they need to choose one among several hypothetical products or behaviors (each characterized by a set of attributes) or the status-quo option to not adopt any [10]. The two experiments cover two choice contexts where the complex contagion theory is expected to apply [8]: (1) political views support [11] (where decision-makers choose which energy policy to support, if any; hereafter referred to as PS experiment) and (2) new social technology adoption [7] (where decision-makers select which instant messaging app to switch to, if any; hereafter referred to as the AA experiment; see Supplementary Note I.1 for all details). In our implementation of the surveys with artificial agents, we prompted the LLMs by including the demographic information and political orientation of the human subjects from ref. [6], which is expected to reduce differences between the human participants and the LLM agents [12] (see Supplementary Note I.2 for all the implementation details). We analyze the resulting choice data in the same way as done by ref. [6] for human subjects, which allows us to estimate the threshold of each artificial agent for each alternative product (see Methods). This procedure places us in the position to directly compare the threshold distribution across human subjects and artificial agents.
In contrast with prior hypotheses [9], we find that the artificial agents exhibit stronger susceptibility to social influence than human subjects. Compared to human subjects’ choices, the social signal is indeed more important in determining artificial agents’ choices in both the policy support and the app adoption experiments. Specifically, for each respondent $n$ and attribute $a$, we measure the difference between the maximum and minimum marginal utility across all the attribute’s levels; We normalize the obtained attribute importance values so that they sum to one for each individual [10]. We find that in the AA experiment, the social signal is the most important attribute for both human and artificial agents, with an average attribute importance of $41.4\%$ (SEM $=0.8\%$) for the human subjects and even higher attribute importance for the artificial agents, $54.0\%$ (SE $=0.4\%$) for GPT and $70.0\%$ (SEM $=0.2\%$) for Gemini, respectively (see Fig. 1D). More surprisingly, in the PS experiment, the social signal is the most important attribute for the artificial agents ($51.4\%$ (SEM $=0.4\%$) for GPT and $32.4\%$ (SEM $=0.2\%$) for Gemini), although it was only ranked 6th out of 7th for the human subjects (average importance $12.8\%$ (SEM $=0.5\%$), see Fig. 1A). With minor discrepancies, the rankings of other attributes’ importance are largely consistent across artificial and human subjects (see Fig. 1A, D).
@@
These findings indicate that the social signal has the highest importance for the artificial agents’ choices even in a context where it was not important for the human subjects’ choices (i.e., in the PS experiment). According to the random utility theory, the threshold of a decision-maker for a given alternative is expected to be inversely proportional to the decision-maker’s susceptibility to social influence [7, 6]. Therefore, we expect the high importance of social signals for the artificial agents’ choices in the PS experiment to directly translate into a lower threshold, which is confirmed by our results (Fig. 1B, E). Specifically, in the PS experiment, the average threshold over all policies for the human participants was $41.2\%$ (SEM $=0.5\%$), while it is only $11.3\%$ (SEM $=0.2\%$)/$12.2\%$ (SEM $=0.2\%$) for the artificial agents generated with GPT/Gemini. Similarly, in the AA experiment, the average threshold for the human participants was $59.6\%$ (SEM $=0.4\%$), which is significantly higher than $1.3\%$ (SEM $=0.1\%$)/$18.0\%$ (SEM $=0.1\%$) observed for the artificial agents generated with GPT/Gemini.
@@
The relevance of the threshold distributions for diffusion outcomes [6] motivates us to examine theoretically whether complex contagions would spread wider in social networks composed of artificial agents compared to networks composed of human subjects. To answer this question, we perform calibrated simulations of complex contagion on empirical networks, where a proportion $1-q$ of the nodes is occupied by human agents, a proportion $q$ by artificial agents (see Fig. 2); the human and artificial agents make adoption decisions determined by their thresholds estimated in the conjoint experiments with human subjects and artificial agents, respectively (see Methods). For each study and for each type of artificial agent, we measure the adoption rate for 36 products/behaviors in 18 empirical networks for 6 values of $q$ under two seeding policies (top degree and random), which leads to a total of 31,104 distinct diffusion simulations (see Supplementary Note II).
Across both contexts, the net effect of $q$ (proportion of artificial agents) is to monotonically widen the reach of the complex contagion. This holds true both in the PS experiment (Fig. 1C) and the AA experiment (Fig. 1F), both when the contagion originates from central nodes and when it originates from randomly-selected nodes. These findings indicate that artificial subjects act as “contagion amplifiers”: For example, in a social network comprising $20\%$ GPT agents with 1% randomly seeded nodes, on average, energy policies would receive 1.14 times more support than in a human-only network; Similarly, under the same conditions, 2.22 times more nodes would decide to switch to a new messaging app compared to a human-only network.
@@ Discussion
These findings open two main directions for future research. One thought-provoking conjecture – supported by the simulated diffusion patterns reported here – is that once artificial agents are injected into a social system, contagion will spread wider. At the same time, because of the limited scope of this Report, it remains open to test the amplifier effect in real-world platforms where artificial and human decision-makers interact longitudinally [13], and to delimit in what conditions we can expect it to occur.
@@
In those contexts where the amplifier effect will be confirmed, we can expect faster behavioral changes at the collective level, with potentially unpredictable consequences. Particularly in high-uncertainty environments, where social signals play a stronger role in guiding behavior [8], artificial agents may accelerate diffusion in ways that are difficult to anticipate. This raises a crucial question for policymakers: how can we steward the resulting collective behavioral patterns toward healthy and sustainable outcomes? Addressing this challenge will likely require integration of experimental and simulation techniques [1, 6], ensuring that artificial agents contribute to beneficial rather than destabilizing social dynamics.
@@
A second important conjecture is that using current LLMs to predict the reach of a social contagion process may systematically overestimate the actual size of contagion. While pioneering studies have highlighted the LLMs’ ability to mirror human behaviors – including cognitive biases [14] and troubling stereotypes [15] – our findings suggest a need for greater scrutiny in contexts where social influence plays a major role. Specifically, future research should investigate the conditions under which LLMs can accurately predict collective human behavior and how their design and prompting can be optimized for more realistic simulations. Advancing this understanding will be crucial for ensuring that AI-driven models of social behavior produce reliable insights rather than misleading projections.
@@
Finally, an important assumption in the analyzed experiments is that all agents – human and artificial alike – are only influenced by the aggregate social signal they receive about the alternatives, encoded in the fraction of friends who have already adopted it. In real-world scenarios, the agents might be influenced differently from agents of the same or different type. Future studies should examine whether knowing that some agents are artificial alter decision-making and diffusion processes. Investigating this question will help refine our understanding of hybrid human-AI social systems and their emergent properties.
Materials and Methods
The threshold estimation follows the methodology described in ref. [6]. Individual $n$’s utility from adopting product or behavior $i$ ($U_{ni}$) consists of two components: the utility of the product attributes on $n$ $(U^{(A)}_{ni})$ and the social influence utility ($U^{(S)}_{ni}$), such that: $U_{ni}=U^{(A)}_{ni}+U^{(S)}_{ni}$. The adoption threshold is then defined as the minimal level of the social signal required for $U_{ni}$ to exceed the status-quo utility ($U^{(0)}_{n}$). Under the assumption that $U_{n}^{(0)}-U^{(A)}_{ni}>0$ and the utility from the social signal is linear in the number of adopters ($U^{(S)}_{ni}=\gamma_{n}s_{ni}$, where $s_{ni}$ denotes the percentage of adopters of $i$ within $n$’s social neighborhood, and $\gamma_{n}$ represents $n$’s marginal utility of the social signal), the adoption threshold can be expressed as in ref. [7, 6]: $\tau_{ni}=(U_{n}^{(0)}-U^{(A)}_{ni})/\gamma_{n}$. Following discrete choice theory [7, 6], the utility is assumed to be additive over product/behavior attributes: $U^{(A)}_{ni}=\sum_{k=1}^{K}\beta_{nk}\,x_{ki}$, where $x_{ki}$ encodes if alternative $i$ has attribute $k$ and $\beta_{nk}$ is the (partworth) utility of $n$ for attribute $k$. All parameters ($\beta_{nk}$, $\gamma_{n}$, $U_{n}^{(0)}$) can be estimated with the Hierarchical Bayes (HB) algorithm from observed choices, as in a choice-based conjoint analysis.
@@ I.1.2 The app adoption study (AA)
In the original study 300 US participants have been recruited from a convenience sample on Prolific. Participants were told to imagine a new instant messaging app available on the market. Participants were presented which 14 choice sets, each consisting of three instant messaging apps and a none option. The apps differed in terms of accessibility (mobile/web), authentication (simple/two-factor/multi-factor), customization level (low/medium/high), support for video calls (multi-person/one-to-one), and the (hypothetical) percentage of friends who are already using the app (1%/23%/45%/76%/98%). Participants had to select from each choice set which app they would use or to select the none option. Each participant received a different set of choice tasks, generated by the Sawtooth software using the balanced overlap method.

### v2

@@ Abstract
Recent advances in artificial intelligence have led to the proliferation of artificial agents in social contexts, ranging from education to online social media and financial markets, among many others. The increasing rate at which artificial and human agents interact makes it urgent to understand the consequences of human-machine interactions for the propagation of new ideas, products, and behaviors in society. Across two distinct empirical contexts, we find here that artificial agents lead to significantly faster and wider social contagion. To this end, we replicate a choice experiment previously conducted with human subjects by using artificial agents powered by large language models (LLMs). We use the experimental results to measure the adoption thresholds of artificial agents and their impact on the spread of social contagion. We find that artificial agents tend to exhibit lower adoption thresholds than humans, which leads to wider network-based social contagions. Our findings suggest that the increased presence of artificial agents in real-world networks may accelerate behavioral shifts, potentially in unforeseen ways.
The increasing role of artificial agents in human activities is not guaranteed to promote more inclusive, healthy, and sustainable societies [1]. As a result, even before the recent popularization of large language models (LLMs), scholars from multiple disciplines have called for the development of an interdisciplinary science of machine behavior where the behavior of AI agents is studied with similar tools to those traditionally used to understand human behaviors [2]. The proliferation of large language models such as ChatGPT and the increased impact of algorithms on large-scale social dynamics makes it urgent to advance in this direction [1, 3]. However, despite some early efforts [4, 5], the impacts of autonomous agents driven by LLMs (hereafter, artificial agents) on social contagion processes have not yet been widely examined. Here we ask: What role could artificial agents play in the spreading of social contagions? To answer, we consider simple artificial agents obtained by prompting popular LLMs (OpenAI’s gpt-3.5-turbo and Google’s gemini-1.5-flash) with the demographic information and political orientation of the human subjects in a previous study [6]. We perform the first measurement of the artificial agents’ threshold, namely, the level of social reinforcement they need before supporting a new policy or adopting a new technology [7]. Our findings indicate that across both contexts, compared to human subjects, artificial agents exhibit a lower threshold, which invariably results in wider diffusion. The higher the proportion of artificial agents in a human-LLM network, the wider the diffusion.
@@
Here, we aim to advance this debate by providing empirical evidence on the artificial agents’ thresholds. To this end, we adopt two different LLMs – OpenAI’s gpt-3.5-turbo model (hereafter, GPT) and Google’s gemini-flash-1.5 model (hereafter, Gemini) – to simulate the responses of survey participants to the same two choice-based conjoint experiments performed in ref. [6] with human subjects. In conjoint experiments, participants face a set of choice tasks where they need to choose one among several hypothetical products or behaviors (each characterized by a set of attributes) or the status-quo option to not adopt any [10]. The two experiments cover two choice contexts where the complex contagion theory is expected to apply [8]: (1) political views support [11] (where decision-makers choose which energy policy to support, if any; hereafter referred to as PS experiment) and (2) new social technology adoption [7] (where decision-makers select which instant messaging app to switch to, if any; hereafter referred to as the AA experiment; see Supplementary Note I.1 for all details). In our implementation of the surveys with artificial agents, we prompted the LLMs by including the demographic information and political orientation of the human subjects from ref. [6], which is expected to reduce differences between human participants and LLM agents [12] (see Supplementary Note I.2 for all the implementation details).
We analyze the resulting choice data through the method adopted by ref. [6] for human subjects, which allows us to estimate the threshold of each artificial agent for each alternative product (see Supplementary Note I.3). Rooted in random utility theory, the method assumes that an individual’s utility from adopting is linear in the social signal received by the individual about the product. Thus, one can derive a simple formula for the individual threshold [7], whose parameters can be estimated via a Hierarchical Bayes algorithm as in standard choice-based conjoint analysis [6]. This procedure places us in an ideal position to directly compare the threshold distribution between human subjects and artificial agents.
@@
In contrast to previous hypotheses [9], we find that artificial agents exhibit a higher susceptibility to social influence than human subjects. In particular, the social signal has high importance (see Supplementary Note I.3 for the importance measurement details) for the artificial agents’ choices even in a context where it was not important for the human subjects’ choices (i.e., in the PS experiment). Specifically in the AA experiment, the social signal is the most important attribute for both human and artificial agents. More surprisingly, in the PS experiment, the social signal is the first and second important attribute for the artificial agents driven by GPT and Gemini, respectively [average attribute importance $51.4\%$ (SEM $=0.4\%$) and $32.4\%$ (SEM $=0.2\%$) for GPT and Gemini, respectively], although it was only ranked 6th out of 7th for human subjects (average importance $12.8\%$ (SEM $=0.5\%$), see Fig. 1A). With minor discrepancies, the rankings of other attributes’ importance are largely consistent across artificial and human subjects (see Figs. 1A, D). From random utility theory (see Supplementary Note I.3), the threshold of a decision maker for a given alternative decreases with their susceptibility to social influence [7, 6]. Therefore, we expect the higher importance of social signals for the artificial agents’ choices in the PS experiment to translate into a lower threshold, which is confirmed by our results (Fig. 1B, E).
@@
The relevance of the threshold distributions for diffusion outcomes [6] motivates us to examine theoretically whether complex contagions would spread wider in social networks composed of artificial agents compared to networks composed of human subjects. To answer this question, we perform calibrated simulations of complex contagion on empirical networks, where a proportion $1-q$ of the nodes is occupied by human agents, a proportion $q$ by artificial agents (see Fig. 2); the human and artificial agents make adoption decisions determined by their thresholds estimated in the conjoint experiments with human subjects and artificial agents, respectively (see Supplementary Note II). For each study and for each LLM, we measure the adoption rate for 36 products/behaviors in 18 empirical networks for 6 values of $q$ under two seeding policies (top degree and random), which leads to 31,104 distinct diffusion simulations (see Supplementary Note II).
Across both contexts, increasing the proportion $q$ of artificial agents widens the reach of the complex contagion. This holds true both in the PS experiment (Fig. 1C) and the AA experiment (Fig. 1F), both when the contagion originates from central nodes and when it originates from randomly-selected nodes. These findings indicate that artificial subjects act as “contagion amplifiers”: For example, in a social network comprising $20\%$ GPT agents with 1% randomly seeded nodes, on average, energy policies would receive 1.14 times more support than in a human-only network; Similarly, under the same conditions, 2.22 times more nodes would decide to switch to a new messaging app compared to a human-only network.
@@ Discussion
Our findings suggest three promising avenues for future research. One thought-provoking conjecture – supported by the simulated diffusion patterns reported here – is that once artificial agents are injected into a social system, contagion will spread wider. At the same time, it remains open to testing the amplifier effect on real-world platforms where artificial and human decision-makers interact longitudinally [13], and to delimit in what conditions we can expect it to occur. Particularly in high-uncertainty environments, where social signals play a stronger role in guiding behavior [8], artificial agents may accelerate diffusion in ways that are difficult to anticipate. This raises a crucial question for policymakers: How can the resulting collective behavioral patterns be stewarded toward healthy systemic outcomes? Addressing this challenge will likely require the integration of experimental and simulation techniques [1, 6], ensuring that artificial agents contribute to beneficial rather than destabilizing social dynamics.
@@
A second important conjecture is that using current LLMs to predict the reach of a social contagion process may systematically overestimate the actual size of the contagion. While pioneering studies have highlighted the LLMs’ ability to mirror human behaviors – including cognitive biases [14] and troubling stereotypes [15] – our findings suggest a need for greater scrutiny in contexts where social influence plays a major role. Future research should investigate the conditions under which LLMs can accurately predict collective human behavior and how their design and prompting can be optimized for more realistic simulations.
@@
Finally, an important assumption in the analyzed experiments is that the agents are only influenced by the aggregate social signal they receive about the alternatives, encoded in the fraction of friends who have already adopted it. Future studies should investigate whether agents’ ability to differentiate between human and artificial counterparts – when such ability is present – influences their decision-making and diffusion processes, thereby refining our understanding of hybrid human-AI systems and their emergent properties.
@@ I.1.2 The app adoption study (AA)
In the original study 300 US participants have been recruited from a convenience sample on Prolific. Participants were told to imagine a new instant messaging app available on the market. Participants were presented with 14 choice sets, each consisting of three instant messaging apps and a none option. The apps differed in terms of accessibility (mobile/web), authentication (simple/two-factor/multi-factor), customization level (low/medium/high), support for video calls (multi-person/one-to-one), and the (hypothetical) percentage of friends who are already using the app (1%/23%/45%/76%/98%). Participants had to select from each choice set which app they would use or to select the none option. Each participant received a different set of choice tasks, generated by the Sawtooth software using the balanced overlap method.
@@ I.2.2 Prompt example AA Study
I.3 Choice data analysis
Threshold estimation.
The threshold estimation follows the methodology described in ref. [6]. Individual $n$’s utility from adopting product or behavior $i$ ($U_{ni}$) is the sum of two terms: the utility of the product attributes on $n$ $(U^{(A)}_{ni})$ and the utility derived from the social signal received by $n$ about $i$ ($U^{(S)}_{ni}$): $U_{ni}=U^{(A)}_{ni}+U^{(S)}_{ni}$. The adoption threshold is then defined as the minimal level of the social signal for which the utility of $n$ from adopting $i$ exceeds $n$’s status-quo utility ($U^{(0)}_{n}$). Under the assumption that $U_{n}^{(0)}-U^{(A)}_{ni}>0$ and the utility from the social signal is linear in the number of adopters ($U^{(S)}_{ni}=\gamma_{n}s_{ni}$, where $s_{ni}$ denotes the percentage of adopters of $i$ within $n$’s social neighborhood, and $\gamma_{n}$ represents $n$’s marginal utility of the social signal), the adoption threshold can be expressed as in ref. [7, 6]: $\tau_{ni}=(U_{n}^{(0)}-U^{(A)}_{ni})/\gamma_{n}$. Following discrete choice theory [7, 6], the utility is assumed to be additive in the product’s or behavior’s attributes: $U^{(A)}_{ni}=\sum_{k=1}^{K}\beta_{nk}\,x_{ki}$, where $x_{ki}$ encodes if alternative $i$ has attribute $k$ and $\beta_{nk}$ is the (partworth) utility of $n$ for attribute $k$. All parameters ($\beta_{nk}$, $\gamma_{n}$, $U_{n}^{(0)}$) can be estimated with the Hierarchical Bayes (HB) algorithm from observed choices, as in a choice-based conjoint analysis [17]. As in ref. [6], we used the R package ChoiceModelR [18] for the HB estimation. The threshold estimation method adopted here has been validated by ref. [6] by showing that (1) it reliably reconstructs the ground-truth thresholds in synthetic data and (2) the individual thresholds it produces can be used to make reliable out-of-sample predictions of individual choices. See ref. [6] for complete details.
Attribute importance measurement.
To measure the importance of attribute $a$ for respondent $n$, we measure the difference between the maximum and minimum marginal utility across all the attribute’s levels; We normalize the obtained attribute importance values so that they sum to one for each individual [10].